undeadknight26 (undeadknight26):

Which of the following equations have the same solution? Solve all three equations separately, and explain each step.

OpenStudy (hugsnotughs):

Bby, I need the equations.

undeadknight26 (undeadknight26):

I. 2x + 4 = 3x – 11 II. 3x + 3 = 51

undeadknight26 (undeadknight26):

2x + 4 = 3x – 11 2x + 3x = 4 - 11 5x = 7 /5 /5 x = 1.4?

OpenStudy (hugsnotughs):

No, I believe you are wrong. \[2x+4=3x-11\]So now, just subtract 2x from both sides, which will get you...\[4=x-11\](remember 1x equals x) Now, you want to add 11 to both sides which will get you...\[15=x\]

undeadknight26 (undeadknight26):

Mind blown :D Thanks!

OpenStudy (hugsnotughs):

A pleasure.

undeadknight26 (undeadknight26):

One more?

OpenStudy (hugsnotughs):

Of course, Mr. Grey.

OpenStudy (hugsnotughs):

This one actually seems really hard, but it is actually pretty simple! Find the LCD for both fraction-equations or whatever the hell you want to call them.

undeadknight26 (undeadknight26):

the lowest common denomiator would be 2 @hugsnkisses ?

undeadknight26 (undeadknight26):

Wrong person @hugsnotughs

OpenStudy (hugsnotughs):

No, the LCD is six. Remember, Least COMMON Denominator. :)

OpenStudy (mathmath333):

ita LCM (least common multiple)

OpenStudy (mathmath333):

LCD-->liquid crystal diode

undeadknight26 (undeadknight26):

*stares* Im confused and my brain is hurting ;-;

OpenStudy (mathmath333):

multiply both sides by 6 and solve further

OpenStudy (hugsnotughs):

^

OpenStudy (hugsnotughs):

The 6 will take out the fractions..

undeadknight26 (undeadknight26):

Can you show me? Im confused ;-;

OpenStudy (mathmath333):